• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Hall A Techs

    Log entry time 16:40:03 on March 25, 2014

    Entry number 387533


    keyword=Lead Flow Adjustment

    Left And Right Dipole lead flows A and B have been adjusted to 75 slm

    A copy of this log entry has been emailed to: andrewl@jlab.org, fansler@jlab.org